# Hybrid Quantum-Classical Recurrent Neural Networks

### v1

@@ Abstract
We present a hybrid quantum-classical recurrent neural network (QRNN) architecture in which the entire recurrent core is realized as a parametrized quantum circuit (PQC) controlled by a classical feedforward network. The hidden state is the quantum state of an $n$-qubit PQC, residing in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$. The PQC is unitary by construction, making the hidden-state evolution norm-preserving without external constraints. At each timestep, mid-circuit readouts are combined with the input embedding and processed by the feedforward network, which provides explicit classical nonlinearity. The outputs parametrize the PQC, which updates the hidden state via unitary dynamics. The QRNN is compact and physically consistent, and it unifies (i) unitary recurrence as a high-capacity memory, (ii) partial observation via mid-circuit measurements, and (iii) nonlinear classical control for input-conditioned parametrization. We evaluate the model in simulation with up to 14 qubits on sentiment analysis, MNIST, permuted MNIST, copying memory, and language modeling, adopting projective measurements as a limiting case to obtain mid-circuit readouts while maintaining a coherent recurrent quantum memory. We further devise a soft attention mechanism over the mid-circuit readouts in a sequence-to-sequence model and show its effectiveness for machine translation. To our knowledge, this is the first model (RNN or otherwise) grounded in quantum operations to achieve competitive performance against strong classical baselines across a broad class of sequence-learning tasks.
@@ 1 Introduction
Figure 1: Hybrid QRNN. (a) Recurrent core PQC with $n=4$ qubits (illustrative) and 16 parametrized gates, acting on a quantum state in the Hilbert space $\mathbb{C}^{2^{n}}$; each horizontal line corresponds to one qubit. (b) $\mathbf{RX}$ and $\mathbf{RY}$ gates.22footnotemark: 2 Each gate in the PQC is parametrized by a rotation angle $\theta_{i}$, where $1\leq i\leq 16$. (c) QRNN unrolled for a sequence of length $k$. The feedforward network $\mathcal{F}$ takes as input the concatenation of the mid-circuit readout vector from the previous timestep and the current input $(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$, and outputs $\bm{\theta}_{t}\in\mathbb{R}^{16}$ containing the 16 rotation angles that parametrize the PQC shown in (a) as $U(\bm{\theta}_{t})$, which acts on the quantum state propagated from the previous step. Refer to caption denotes qubit measurements.
In this work, we present a hybrid quantum–classical recurrent neural network (QRNN) architecture grounded in quantum operations, in which the entire recurrent core is realized as a PQC. The hidden state is the quantum state of the PQC, residing in an exponentially large Hilbert space. A classical feedforward network parametrizes and steers the quantum computation, introducing nonlinearity through mid-circuit readouts, which are realized in simulation as a limiting case via projective measurements. This avoids emulating nonlinearity through linear quantum dynamics, leaving the PQC strictly for coherent unitary evolution.
Fig. 2 illustrates both the PQC (with four qubits shown for illustration) and the unrolled QRNN:
  • At each timestep $t$, the input is mapped to a classical embedding $\mathbf{x}_{t}$ via a learnable embedding layer.
  • A classical feedforward network $\mathcal{F}$ takes as input the concatenation of the readout vector $\mathbf{z}_{t-1}$ (outputs from all measurements at timestep t1t-1) and the current input $\mathbf{x}_{t}$. It outputs the PQC parameters $\bm{\theta}_{t}$ which configure the PQC with a fixed gate layout (Fig. 1(a)), denoted $U(\bm{\theta}_{t})$, applied at timestep $t$ (Fig. 1(c)).
@@
  • The mid-circuit readout $\mathbf{z}_{t}$ (or the final readout at the end of the sequence) is a classical feature vector obtained from the quantum state via measurements and is used: (i) as recurrent feedback $\mathbf{z}_{t-1}$ at timestep $t$, and (ii) as the input to task-specific classical layers.
We develop the models on GPUs, allowing us to simulate and train quantum recurrence via classical backpropagation, with the expectation that such models will become classically unsimulatable as the number of qubits increases. To our knowledge, this is the first model grounded in quantum operations (RNN or otherwise) demonstrated in classical simulation with up to 14 qubits across six realistic sequence-modeling tasks, achieving competitive performance with LSTM and the scaled Cayley orthogonal scoRNN designed for norm preservation [Helfrich et al., 2018]. Experiments also show that classical nonlinear control and feedback are effective, with the nonlinear variants outperforming their linear counterparts, and that the unitary quantum recurrent core maintains more stable gradients than LSTMs (§4.6).
@@ 2 Related Work
Bausch [Bausch, 2020] developed a QRNN with a persistent quantum memory based on quantum neurons [Cao et al., 2017]. However, nonlinearities are emulated within PQCs using probabilistic repeat-until-success circuit execution and postselection steps. The inherently linear nature of quantum computation constrains such emulation, and the available forms of nonlinearity remain limited [Yan et al., 2020, Moreira et al., 2023, Zi et al., 2024].
@@
Li et al. [2023] and Siemaszko et al. [2023] also model recurrences with PQCs, while supporting per-timestep readouts, but they rely entirely on linear quantum dynamics of the PQC without explicit nonlinearities or classical control.
@@ 3.1 PQC
Measurements.
To probe the quantum state of a PQC, we can perform measurements to obtain real-valued readouts. These readouts provide partial observations of the state, and any required number of measurements, on any of the qubits, can be combined for downstream tasks. For instance, a measurement through the Pauli-$\mathbf{Z}$ observable with the unitary
$\mathbf{Z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
assigns scalar values (e.g., +1+1 for $|0\rangle$ and 1-1 for $|1\rangle$) in the computational basis in the single-qubit case. For a general quantum state however, the outcome is probabilistic: it yields +1+1 with probability $|\alpha|^{2}$ and 1-1 with probability $|\beta|^{2}$. The expectation value of this measurement is given by $|\alpha|^{2}-|\beta|^{2}$, which can be used as a real-valued readout in hybrid quantum-classical models.
Although the readouts obtained via measurement are a nonlinear function of the gate parameters, particularly those used in parametrized rotation gates such as $\mathbf{RX}$, the resulting nonlinearity is generally weak (§4).
@@ 3.2 Hybrid Model
Let $\mathbf{x}_{t}$ be the input embedding at timestep $t$, and let $\mathbf{z}_{t-1}$ be the measurement-based readout from the previous timestep. In the most generic form of the hybrid model, [^7] We may add extra transformations to the measurement outcomes before classifications or feeding them to the next step; see §4. the two are concatenated into a single vector $\mathbf{u}_{t}=(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$ and passed through a classical feedforward network $\mathcal{F}$ with one hidden layer and a nonlinearity.
@@
The PQC itself is defined by a unitary operator $U(\bm{\theta}_{t})$, parametrized by $\bm{\theta}_{t}$. [^8] We use $U(\bm{\theta}_{t})$ to denote all unitary operations composed of multiple parametrized gates, each acting on one or more qubits with parameters drawn from $\bm{\theta}_{t}$. Applying the gates in $U(\bm{\theta}_{t})$ to the quantum state $\mathbf{h}_{t-1}=|\psi_{t-1}\rangle$ yields the updated state $\mathbf{h}_{t}=U(\bm{\theta}_{t})\,|\psi_{t-1}\rangle$. The updated state is then measured to obtain a classical readout vector
$\mathbf{z}_{t}=\textbf{Measure}(\mathbf{h}_{t}),$ (3)
which serves as a proxy for the quantum state and is combined with the next input $\mathbf{x}_{t+1}$ to evolve the recurrence. Mid-circuit measurements are simulated without collapsing the quantum state via projective measurements, thereby preserving coherence across timesteps.
We train the entire hybrid model end-to-end using classical backpropagation, optimizing the parameters $\bm{\Theta}=\{\mathbf{W}_{1},\mathbf{b}_{1},\mathbf{W}_{2},\mathbf{b}_{2}\}$ via standard optimizers, such as Adam [Kingma and Ba, 2014]. Because each $\mathbf{z}_{t}$ is real-valued, it can be used both as a per-timestep output and as a contextual embedding for soft attention in sequence-to-sequence decoding.
@@ 4 Experiments
For $\mathbf{Measure}$ in Eq. 3, measurements are performed in each of the $\mathbf{X}, Pauli$-$\mathbf{Y}$, and Pauli-$\mathbf{Z}$ observables across all wires in the PQC, and the measurement outcomes are combined to form $\mathbf{z}_{t}$ (Eq. 3). For the feedforward network $\mathcal{F}$ (Eq. 1 and Eq. 2), we experimented with ReLU, LeakyReLU, GLU and GELU nonlinearities. [^10] GLU requires projecting to twice the output dimensionality, effectively increasing the parameter count compared to standard nonlinearities like ReLU, when all other dimensions are held constant. For both language modeling and translation, we first transform the measurement outcomes with a separate feedforward layer and use the result both for vocabulary classification and as input to the next timestep.
@@ 4.1 Sentiment Analysis
Table 1: Classification accuracy on IMDB. Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp. $\dagger$ indicates the LSTM in Dai and Le [2015].
@@
The hybrid model for this task follows the generic hybrid architecture described in §3.2. At the final input token, we apply an affine transformation to the measurement outcomes to produce two logits, which are used for classification via cross-entropy. Table 1 summarizes the results. QRNNLeakyReLU achieves the highest test accuracy. Ablating the classical nonlinearity (Eq. 1) degrades performance, though increasing the number of qubits in the linear model still yields some accuracy gains. Adding the nonlinearity results in a substantial improvement, outperforming all baselines. On this task, the orthogonal scoRNN underperforms other models, despite having a larger hidden state and over five times more parameters.
4.2 MNIST and Permuted-MNIST
Table 2: Classification accuracy on MNIST and pMNIST. Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp. $\dagger$ indicates the QRNN model of [Bausch, 2020] with 13 qubits and each digit downscaled to $4\times 4$ and binarized.
@@ 4.4 Word-Level Language Modeling
Table 3: PTB word-level language modeling (PPL). Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
@@ 4.5 Machine Translation
Table 4: Multi30K German-to-English translation (BLEU). Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
@@
Results in Table 4 show that QRNNGLU with 13 qubits closely matches the LSTM, followed by QRNNGLU with 10 qubits. For the QRNN, it is somewhat surprising that intermediate readouts can still support mechanisms like soft attention, since these readouts capture only partial projections of the quantum state rather than the full hidden state. This suggests that, despite intermediate measurements, sufficient information is retained and propagated across timesteps. We qualitatively interpret the learned soft alignments on a few examples where the translations required non-trivial linguistic interpretations in Appendix D.
@@ 5 Discussion and Conclusion
Different quantum hardware platforms currently require distinct control stacks, and architectural choices do not translate one-to-one across devices, with factors such as native gate sets, qubit connectivity, and the implementation of mid-circuit measurements all affecting the realization of a given circuit. The aim here is not to prescribe a hardware roadmap but to analyze a hardware-realistic base case under idealized classical simulation to study the empirical properties of the architecture, where we model mid-circuit observations via projective measurements as a limiting case.
@@
This paper bridges quantum operations and recurrent learning by introducing a new hybrid QRNN whose recurrent core is implemented as a PQC steered by a classical controller. The unitary dynamics preserve norms, promoting stable gradient propagation; the controller supplies the nonlinearity and task adaptivity needed for expressiveness; and mid-circuit measurements enable per-timestep readouts and classical feedback. As techniques improve [Abbas et al., 2023] and quantum hardware matures, the architecture provides a path toward hardware-realistic quantum models for sequential learning.

### v2

@@ Abstract
We present a hybrid quantum-classical recurrent neural network (QRNN) architecture in which the recurrent core is realized as a parametrized quantum circuit (PQC) controlled by a classical feedforward network. The hidden state is the quantum state of an $n$-qubit PQC in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$, which serves as a coherent recurrent quantum memory. The PQC is unitary by construction, making the hidden-state evolution norm-preserving without external constraints. At each timestep, mid-circuit Pauli expectation-value readouts are combined with the input embedding and processed by the feedforward network, which provides explicit classical nonlinearity. The outputs parametrize the PQC, which updates the hidden state via unitary dynamics. The QRNN is compact and physically consistent, and it unifies (i) unitary recurrence as a high-capacity memory, (ii) partial observation via mid-circuit readouts, and (iii) nonlinear classical control for input-conditioned parametrization. We evaluate the model in simulation with up to 14 qubits on sentiment analysis, MNIST, permuted MNIST, copying memory, and language modeling. For sequence-to-sequence learning, we further devise a soft attention mechanism over the mid-circuit readouts and show its effectiveness for machine translation. To our knowledge, this is the first model (RNN or otherwise) grounded in quantum operations to achieve competitive performance against strong classical baselines across a broad class of sequence-learning tasks.
@@ 1 Introduction
Figure 1: Hybrid QRNN. (a) Recurrent core PQC with $n=4$ qubits (illustrative) and 16 parametrized gates, acting on a quantum state in the Hilbert space $\mathbb{C}^{2^{n}}$; each horizontal line corresponds to one qubit. (b) $\mathbf{RX}$ and $\mathbf{RY}$ gates.22footnotemark: 2 Each gate in the PQC is parametrized by a rotation angle $\theta_{i}$, where $1\leq i\leq 16$. (c) QRNN unrolled for a sequence of length $k$. The feedforward network $\mathcal{F}$ takes as input the concatenation of the mid-circuit readout vector from the previous timestep and the current input $(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$, and outputs $\bm{\theta}_{t}\in\mathbb{R}^{16}$ containing the 16 rotation angles that parametrize the PQC shown in (a) as $U(\bm{\theta}_{t})$, which acts on the quantum state propagated from the previous step. Refer to caption denotes qubit expectation-value readouts.
In this work, we present a hybrid quantum-classical QRNN in which the recurrent core is realized as a PQC controlled by a classical feedforward network. The hidden state is the quantum state of the $n$-qubit PQC, residing in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$, which provides a high-capacity coherent quantum memory. Nonlinearity is introduced through classical computation rather than approximated within the quantum circuit, leaving the PQC strictly for unitary evolution of the hidden state. Fig. 2 illustrates both the PQC (with four qubits shown for illustration) and the unrolled QRNN:
  • At timestep $t$, the input is mapped to a learnable representation $\mathbf{x}_{t}$ via an embedding layer.
  • A classical feedforward network $\mathcal{F}$ takes as input the concatenation of $\mathbf{z}_{t-1}$ (readout outputs at timestep t1t-1) and the current input $\mathbf{x}_{t}$. It outputs the PQC parameters $\bm{\theta}_{t}$ which configure the PQC with a fixed gate layout (Fig. 1(a)), denoted $U(\bm{\theta}_{t})$, applied at timestep $t$ (Fig. 1(c)).
@@
  • The mid-circuit readout $\mathbf{z}_{t}$ (or the final readout at the end of the sequence) is a real-valued vector obtained from the quantum state via Pauli expectation-value readouts and is used: (i) as recurrent feedback $\mathbf{z}_{t-1}$ at timestep $t$, and (ii) as the input to task-specific classical layers.
We develop the models on GPUs, allowing us to simulate and train quantum recurrence via classical backpropagation, with the expectation that such models will become classically unsimulatable as the number of qubits increases. To our knowledge, this is the first model (RNN or otherwise) grounded in quantum operations demonstrated in classical simulation with up to 14 qubits across six realistic sequence-modeling tasks, achieving competitive performance with LSTM and the scaled Cayley orthogonal scoRNN designed for norm preservation [Helfrich et al., 2018]. Experiments also show that classical nonlinear control and feedback are effective, with nonlinear variants outperforming their linear counterparts, and that the unitary quantum recurrent core maintains more stable gradients than LSTMs on sequences of up to 400 tokens (§4.6).
@@ 2 Related Work
Bausch [2020] proposes a QRNN whose recurrence is implemented by iterating a quantum cell built from “quantum neurons”. Nonlinearity is induced inside the circuit via measurement-and-postselection primitives (repeat-until-success), together with amplitude amplification to make the procedure near-deterministic. Consequently, the per-timestep update is not strictly unitary and introduces measurement back-action on the memory. Because the dynamics remain linear between measurements, the effective nonlinear maps achievable by such measurement-based schemes are structurally constrained, and the repertoire of admissible nonlinearities remains limited [Yan et al., 2020, Moreira et al., 2023, Zi et al., 2024].
@@
Li et al. [2023] and Siemaszko et al. [2023] also model recurrences with PQCs and support per-timestep readouts, but they rely entirely on linear quantum dynamics without other nonlinearity or classical control.
@@ 3.1 PQC
Readouts via Pauli expectations.
A succinct way to summarize a quantum state is via expectation values of Pauli observables. For each qubit $k$, the Pauli operators $X_{k},Y_{k},Z_{k}$ are the three fundamental single-qubit observables that act on that qubit (with identity on all other qubits). For example, the Pauli-$Z$ operator measures the computational basis with
$Z\;=\;\begin{pmatrix}1&0\\[2.0pt] 0&-1\end{pmatrix}$,
while $X$ and $Y$ measure superposition states in orthogonal bases. For a single qubit $\lvert\psi\rangle=\alpha\lvert 0\rangle+\beta\lvert 1\rangle$, the Pauli-$Z$ expectation is $\langle Z\rangle=\langle\psi\lvert Z\rvert\psi\rangle=|\alpha|^{2}-|\beta|^{2}\in[-1,1]$. For an $n$-qubit state, per-qubit expectations $\{\langle X_{k}\rangle,\langle Y_{k}\rangle,\langle Z_{k}\rangle\}_{k=1}^{n}$ and correlation terms (Pauli strings) $\langle P_{1}\otimes\cdots\otimes P_{n}\rangle$ with $P_{\ell}\in\{I,X,Y,Z\}$ yield bounded, hardware-agnostic real-valued summaries useful for analysis and downstream modeling.
Although expectation-value readouts are nonlinear functions of parametrized rotation angles (e.g., in $\mathbf{RX}$ gates), the resulting nonlinearity is generally insufficient on its own (§4).
@@ 3.2 Hybrid Model
Let $\mathbf{x}_{t}$ be the input embedding at timestep $t$, and let $\mathbf{z}_{t-1}$ be the readout vector from the previous timestep. In the most generic form of the hybrid model, [^7] Extra transformations may be applied to the readouts before classifications or feeding them to the next step (for some tasks); see §4. the two are concatenated into a single vector $\mathbf{u}_{t}=(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$ and passed through a classical feedforward network $\mathcal{F}$ with one hidden layer and a nonlinearity.
@@
The PQC itself is defined by a unitary operator $U(\bm{\theta}_{t})$ parametrized by $\bm{\theta}_{t}$. [^8] Here $U(\bm{\theta}_{t})$ denotes the product of the circuit’s parametrized gates, each acting on one or more qubits with its parameter drawn from $\bm{\theta}_{t}$. Applying $U(\bm{\theta}_{t})$ to the prior state $\mathbf{h}_{t-1}=\lvert\psi_{t-1}\rangle$ yields the updated state
$\mathbf{h}_{t}=U(\bm{\theta}_{t})\,\lvert\psi_{t-1}\rangle.$
A classical readout vector is then computed from Pauli expectation values:
$\mathbf{z}_{t}=\mathrm{Readout}(\mathbf{h}_{t})=\big(\,\langle X_{1}\rangle_{t},\ \langle Y_{1}\rangle_{t},\ \langle Z_{1}\rangle_{t},\ \ldots,\ \langle X_{n}\rangle_{t},\ \langle Y_{n}\rangle_{t},\ \langle Z_{n}\rangle_{t}\,\big),$ (3)
where $\langle P_{k}\rangle_{t}=\langle\psi_{t}\lvert P_{k}\rvert\psi_{t}\rangle$ for $P_{k}\in\{X_{k},Y_{k},Z_{k}\}$ and $\lvert\psi_{t}\rangle=\mathbf{h}_{t}$. [^9] For computational efficiency, we use only single-qubit expectations rather than multi-qubit Pauli strings.
Serving as a proxy for the quantum state, $\mathbf{z}_{t}$ is concatenated with the next input $\mathbf{x}_{t+1}$ to produce $\bm{\theta}_{t+1}$ via the classical controller. Because the Pauli expectation-value readouts do not alter the state, coherence of the quantum state is preserved across timesteps, allowing it to function as a coherent recurrent memory.
We train the entire hybrid model end-to-end using classical backpropagation, optimizing the parameters $\bm{\Theta}=\{\mathbf{W}_{1},\mathbf{b}_{1},\mathbf{W}_{2},\mathbf{b}_{2}\}$ via standard optimizers, such as Adam [Kingma and Ba, 2014]. For sequence-to-sequence learning, $\mathbf{z}_{t}$ provides per-timestep outputs and serves as contextual embeddings for soft attention decoding.
@@ 4 Experiments
For the feedforward network $\mathcal{F}$ (Eq. 1 and Eq. 2), we experimented with ReLU, LeakyReLU, GLU and GELU nonlinearities. [^11] GLU requires projecting to twice the output dimensionality, effectively increasing the parameter count compared to standard nonlinearities like ReLU, when all other dimensions are held constant. For both language modeling and translation, we first transform the readouts with a separate feedforward layer and use the result both for vocabulary classification and as input to the next timestep.
@@ 4.1 Sentiment Analysis
Table 1: Classification accuracy on IMDB. Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp. $\dagger$ indicates the LSTM in Dai and Le [2015].
@@
The hybrid model for this task follows the generic hybrid architecture described in §3.2. At the final input token, we apply an affine transformation to the readouts to produce two logits, which are used for classification via cross-entropy. Table 1 summarizes the results. QRNNLeakyReLU achieves the highest test accuracy. Ablating the classical nonlinearity (Eq. 1) degrades performance, though increasing the number of qubits in the linear model still yields some accuracy gains. Adding the nonlinearity results in a substantial improvement, outperforming all baselines. On this task, the orthogonal scoRNN underperforms other models, despite having a larger hidden state and over five times more parameters.
4.2 MNIST and Permuted MNIST
Table 2: Classification accuracy on MNIST and pMNIST. Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp. $\dagger$ indicates the QRNN model of [Bausch, 2020] with 13 qubits and each digit downscaled to $4\times 4$ and binarized.
@@ 4.4 Word-Level Language Modeling
Table 3: PTB word-level language modeling (PPL). Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
@@ 4.5 Machine Translation
Table 4: Multi30K German-to-English translation (BLEU). Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
@@
Results in Table 4 show that QRNNGLU with 13 qubits closely matches the LSTM, followed by QRNNGLU with 10 qubits. For the QRNN, it is somewhat surprising that intermediate readouts can still support mechanisms like soft attention, since these readouts capture only partial projections of the quantum state rather than the full hidden state. This suggests that, despite mid-circuit readouts, sufficient information is retained and propagated across timesteps. We qualitatively interpret the learned soft alignments on a few examples where the translations required non-trivial linguistic interpretations in Appendix D.
@@ 5 Discussion and Conclusion
Different quantum hardware platforms currently require distinct control stacks, and architectural choices do not translate one-to-one across devices, with factors such as native gate sets, qubit connectivity, and the implementation of mid-circuit readout all affecting the realization of a given circuit. The aim here is not to prescribe a hardware roadmap but to analyze a hardware-realistic base case under idealized classical simulation to study the empirical properties of the architecture, where we model mid-circuit observations via expectation-value readouts.
@@
This paper bridges quantum operations and recurrent learning by introducing a new hybrid QRNN whose recurrent core is implemented as a PQC steered by a classical controller. The unitary dynamics preserve norms, promoting stable gradient propagation; mid-circuit, per-timestep readouts inject task adaptability; and the classical controller supplies the nonlinearity and feedback for expressiveness. As techniques improve [Abbas et al., 2023] and quantum hardware matures, the architecture provides a path toward hardware-realistic quantum models for sequential learning.